two loading brands of potato chips were tested to compare their sodium content. ten packs of brand a yielded a mean sodium content of 80 mg/pack with a standard deviation of 5.4 mg while eight packs of brand b yielded an average content of 75 mg/pack with standard deviation of 6.1 mg. is there difference in the sodium content of the two brands at the 0.05 level of significance? (assume normality with equal variances)


1
Expert's answer
2021-09-24T01:06:01-0400

Let the following notations denote the 2 populations,

Population 1

n1=10

"\bar{x}"1=80

s1=5.4

Population 2

n2=8

"\bar{x}"2 =75

s2=6.1

The hypothesis tested is,

H0:"\mu"1="\mu"2 against H1:"\mu"1 "\not=" "\mu"2 at "\alpha" =0.05

Assuming normality and equal variances, this hypothesis is tested using the student's t distribution.

Since population variances are unknown and assumed equal, pooled sample variance is used. Its formula is given by,

"Sp^2=((n1-1)s1^2+(n2-1)s2^2)\/(n1+n2-2)"

=((10-1)*5.42+(8-1)*6.12)/(10+8-2)

=9*(29.16)+7*(37.21)/16

=32.681875

The test statistic t*(calculated) is given by,

t*=( "\bar{x}"1- "\bar{x}"2)"\/""\sqrt{Sp^2*(1/n1+1/n2)}"

="(80-75)\/" "\sqrt{32.681875*(1/10+1/8)}"

=5/2.71171936

=1.8438(4 decimal places)

t* is compared with the table value with n1+n2-2=16 degrees of freedom and "\alpha/2=0.025" since it is a two sided t-test.

t0.025, 16=2.120.

Considering that t*(calculated)=1.8438 is less than the table value t0.025, 16=2.120, we fail to reject the null hypothesis and conclude that there is a significant evidence at the 95% level of confidence that sodium content for the two brands are not different.